# Towards Efficient Training of Graph Neural Networks: A Multiscale Approach

### v3

@@ 7 Conclusions
In this work, we propose an efficient approach for training graph neural networks and evaluate its effectiveness across multiple datasets, architectures, and pooling strategies. We demonstrate that all proposed approaches, Coarse-to-Fine, Sub-to-Full, and Multiscale Gradient Computation, improve performance while reducing the training costs.
Our method is adaptable to any architecture, dataset, or pooling technique, making it a versatile tool for GNN training. Our experiments reveal that no single setting is universally optimal; performance varies depending on the dataset and architecture. However, our approach consistently yields improvements across all considered settings.
Appendix A Theorem 4.1 Proof
@@
${\ell}({\boldsymbol{\Theta}})={\frac{1}{2N}}\|{\bf A}{\bf X}{\boldsymbol{\Theta}}-{\bf Y}\|^{2},$ (15)
and the sketching problem from Woodruff (2014)
${\ell}_{C}({\boldsymbol{\Theta}})={\frac{1}{2N}}\|{\bf P}^{\top}{\bf A}{\bf X}{\boldsymbol{\Theta}}-{\bf P}^{\top}{\bf Y}\|^{2}$ (16)
@@
Theorem A.1 (Woodruff (2014), Thm. 23).
Let ${\bf P}^{\top}$ be a sparse subspace embedding with $O(\frac{d^{2}}{\epsilon})$ rows, and let ${\boldsymbol{\Theta}}^{*}$ be the solution of Equation (15) and ${\boldsymbol{\Theta}}_{C}^{*}$ be the solution of Equation (16). Then with probability 0.99, it holds that:
${\ell}_{c}({\boldsymbol{\Theta}}_{C}^{*})={\frac{1}{2N}}\|{\bf P}^{\top}{\bf A}{\bf X}{\boldsymbol{\Theta}}_{C}^{*}-{\bf P}^{\top}{\bf Y}\|^{2}={\frac{1\pm\epsilon}{2N}}\|{\bf A}{\bf X}{\boldsymbol{\Theta}}^{*}-{\bf Y}\|^{2}$ (17)
@@
${\ell}_{c}({\boldsymbol{\Theta}}_{C}^{*})$ $=$ ${\frac{1}{2N}}\|{\bf P}^{\top}{\bf A}{\bf P}{\bf P}^{\top}{\bf X}{\boldsymbol{\Theta}}_{C}^{*}-{\bf P}^{\top}{\bf Y}\|^{2}={\frac{1}{2N}}\|{\bf P}^{\top}{\bf A}{\bf X}{\boldsymbol{\Theta}}_{C}^{*}-{\bf P}^{\top}{\bf Y}+{\bf R}\|^{2}$
@@
Here we used norm properties and Theorem A.1.
@@
Let ${\boldsymbol{\Theta}}^{*}$ be the solution of equation (15) and let ${\boldsymbol{\Theta}}_{C}^{*}$ be the solution of equation (16). Let ${\bf P}^{\top}$ be a sparse subspace embedding with $O(\frac{c^{2}}{\epsilon})$ rows, where $c$ is the size of the feature vectors (number of graph channels), that injects any vector into a subset of indices, which we denote by $C$. According to this subset, denote ${\bf A}=\begin{bmatrix}{\bf A}_{CC}&{\bf A}_{CF}\\ {\bf A}_{FC}&{\bf A}_{FF}\end{bmatrix}$, and ${\bf X}=\begin{bmatrix}{\bf X}_{C}\\ {\bf X}_{F}\end{bmatrix}$. Then if $\sigma=Id$, there exists $\|{\bf R}\|<\|{\bf A}_{CF}{\bf X}_{F}{\boldsymbol{\Theta}}_{C}^{*}\|$ such that with probability .99:
${\ell}_{c}({\boldsymbol{\Theta}}_{C}^{*})={\frac{1}{2N}}\|{\bf P}^{\top}{\bf A}{\bf P}{\bf P}^{\top}{\bf X}{\boldsymbol{\Theta}}_{C}^{*}-{\bf P}^{\top}{\bf Y}\|^{2}\leq{\frac{1\pm\epsilon}{2N}}\left(\|{\bf A}{\bf X}{\boldsymbol{\Theta}}^{*}-{\bf Y}\|^{2}+\|{\bf R}\|^{2}\right)$ (19)
The approximation quality depends on the residual matrix ${\bf R}$. For separable graphs, there are no edges that are omitted and therefore ${\bf R}=0$. As the number of edges that are omitted increases, the solution of the coarse-scale problem diverges from the solution of the fine-scale one.
@@ Appendix B Qtips and MNIST Datasets
We evaluate our methods on the Qtips dataset by constructing a training set of 100 rods and a test set of 1,000 rods. We employ a 4-layer GCN with 256 hidden channels and compare the performance of our approach across three levels of coarsening. The baseline is trained for 500 epochs, while the multiscale methods are trained for [400, 200, 100] epochs across levels. Results, shown in Table 6, indicate that the combination of random coarsening and subgraph pooling achieves the best performance, particularly when using $p=3$.
We further test our approach on a graph dataset derived from the MNIST image dataset LeCun et al. (1998). The original MNIST images consist of $28\times 28$ grayscale pixels, each depicting a digit from 0 to 9. We convert each image into a graph by treating each pixel as a node and connecting nodes using their spatial positions and a $k$-nearest-neighbor algorithm. The task is to assign each node either to a background class or to the digit class it belongs to. We train a 4-layer GCN with 256 hidden channels on 1,000 graphs and evaluate on 100 test graphs. Results are presented in Table 7, where we again observe that the combination of random coarsening and subgraph pooling achieves the best performance among the methods evaluated.
@@
Appendix C Loss Comparison between Multilevel Training and Single-Level Training
To further demonstrate the efficiency of our methods, we show in Figure 5 a comparison between the loss during training, both using single-level training and multiscale training. We show that both training approaches converge to similar loss after a similar number of training epochs, while as demonstrated in Table 2, the cost of a single iteration on the coarse graph is lower than on the original fine graph. The plot shows results for the OGBN-Arxiv dataset Hu et al. (2020), using GCN Kipf & Welling (2016) and random pooling.
@@ Appendix C Loss Comparison between Multilevel Training and Single-Level Training
Appendix D Cora, CiteSeer and PubMed Datasets
We evaluate our method on the Cora, Citeseer, and PubMed datasets Sen et al. (2008) using the standard split from Yang et al. (2016), with 20 nodes per class for training, 500 validation nodes, and 1,000 testing nodes. Our experiments cover standard architectures—GCN Kipf & Welling (2016), GIN Xu et al. (2018), and GAT Velickovic et al. (2017)—and are summarized in Table 8, where we achieve performance that is either superior to or on par with the fine-grid baseline, which trains directly on the original graphs.
We explore two pooling strategies—random and Top-k—along with subgraph generation via $k$-hop neighborhoods and a hybrid approach combining both. Each method is tested under two connectivity settings: the original adjacency ($p=1$) and an enhanced version ($p=2$), where we square the adjacency matrix. We report test accuracy across three runs with different random seeds and include standard deviation to assess robustness. Our results indicate that enhanced connectivity improves accuracy in some settings. Table 8 further presents three levels of multiscale training alongside theoretical FLOP analyses, demonstrating the efficiency of our approach. The details of the FLOPs analysis are provided in Section 5, and additional results for two and four levels of multiscale training are provided in Table 9 and Table 10. We provide further experimental details in Table 16, and additional information about the datasets appears in Table 12.
@@ Appendix D Cora, CiteSeer and PubMed Datasets
  Random(p=2) $2.63\times 10^{2}$ 83.0% (0.5%) $6.82\times 10^{2}$ 66.0% (0.8%) $9.99\times 10^{2}$ 73.8% (0.4%)
@@
  Random(p=1) $2.65\times 10^{2}$ 80.3% (0.2%) $8.11\times 10^{2}$ 72.9% (1.5%) $7.29\times 10^{2}$ 77.5% (1.5%)
@@
Appendix E Additional Transductive Learning datasets
We further test our methods using additional datasets. We provide results for Flickr Zeng et al. (2019) and PPI (transductive) Yang et al. (2020), showing both timing analysis, in Table 11, and full multiscale training results in Table 13. Our results show that while achieving comparable accuracy to the original fine grid, our methods show significant timing improvement both in a theoretical manner as well as in FLOPs analysis, which we calculate as explained in Section 5.
In addition, we test our methods and show results for Facebook, BlogCatalog (Yang et al., 2020), DBLP (Bojchevski & Günnemanng, 2017) and WikiCS (Mernyei & Cangea, 2020), in Table 14 and Table 15. This further illustrates that our methods are not restricted to a specific dataset or network architecture, as our results are consistently comparable to the original training baseline results. Details of our experiments are provided in Table 16, and we list additional details of the datasets in Table 12.
@@ Appendix E Additional Transductive Learning datasets
Subgraphs $4.92\times 10^{3}$ 73.16 54.61% (0.8%) $1.61\times 10^{3}$ 113.90 71.63% (0.3%)
@@
Topk $8.93\times 10^{3}$ 34.47 51.99% (5.1%) $3.54\times 10^{3}$ 33.98 84.92% (0.8%)
@@
Cora Sen et al. (2008) 2,708 10,556 1,433 7
Citeseer Sen et al. (2008) 3,327 9,104 3,703 6
Pubmed Sen et al. (2008) 19,717 88,648 500 3
Flickr Zeng et al. (2019) 89,250 899,756 500 7
WikiCS Mernyei & Cangea (2020) 11,701 216,123 300 10
DBLP Bojchevski & Günnemanng (2017) 17,716 105,734 1,639 4
Facebook Yang et al. (2020) 4,039 88,234 1,283 193
BlogCatalog Yang et al. (2020) 5,196 343,486 8,189 6
PPI (transductive) Yang et al. (2020) 56,944 1,612,348 50 121
@@
Network architecture GCN Kipf & Welling (2016) with 4 layers and 192 hidden channels. GIN Xu et al. (2018) with 3 layers and 256 hidden channels. GAT Velickovic et al. (2017) with 3 layers and 64 hidden channels, using 2 heads
Loss function Negative Log Likelihood Loss except for PPI (transductive) and Facebook Yang et al. (2020) which use Cross Entropy Loss
Optimizer Adam Kingma & Ba (2014)
@@
Sub-to-Full strategy Using ego-networks Gupta et al. (2014). 6 hops on level 2, 4 on level 3, and 2 on level 4
@@
Appendix F Additional Results for Shapenet Dataset
We test out methods on the full Shapenet dataset Chang et al. (2015) using DGCNN Wang et al. (2018) and demonstrate the results of our multiscale training process in Table 17. We observe that our method is not limited to the ’Airplane’ category of this dataset, and we achieved comparable results across various connectivity values and pooling methods while using a significantly cheaper training process.
To emphasize the efficiency of our methods, we further provide timing analysis of Shapanet ’Airplanes’ data computations. Timing was calculated using GPU computations, and we averaged across 5 epochs after the training loss was stabilized. Results appear in Table 18 for multiscale training and in Table 19 for multiscale gradient computations, where we show that our training process was indeed faster than the original process, which only took into consideration the fine grid. We provide additional experimental details in Table 20.

### v5

@@ 8 Conclusions
In this work, we propose an efficient approach for training GNNs and evaluate its effectiveness across multiple datasets, architectures, and pooling strategies. We demonstrate that all proposed approaches, Coarse-to-Fine, Sub-to-Full, and Multiscale Gradient Computation, improve performance while reducing the training costs.
Our method is adaptable to any architecture, dataset, or pooling technique, making it a versatile tool for GNN training. Our experiments reveal that no single setting is universally optimal; performance varies depending on the dataset and architecture. However, our approach consistently yields improvements across all considered settings – for each dataset and network, at least one coarsening method led to improved performance, while the differences between methods were generally minor in practice.
9 Acknowledgements
This research was supported by the Ministry of Innovation, Science and Technology, Israel, and the Ministry of Education and Science of Ukraine as the Joint Ukrainian-Israeli research project (under contract M/38-2024, No. 0124U003512). Also, this research was partially supported by Grant No. 2023771 from the United States-Israel Binational Science Foundation (BSF) and by Grant No. 2411264 from the United States National Science Foundation (NSF).
Appendix A Theorem˜4.1 Proof
@@ 9 Acknowledgements
$\mathcal{L}({\boldsymbol{\theta}})={\frac{1}{2N}}\|{\bf A}{\bf X}{\boldsymbol{\theta}}-{\bf Y}\|^{2},$ (15)
and the sketching problem from woodruff2014sketching
$\mathcal{L}_{C}({\boldsymbol{\theta}})={\frac{1}{2N}}\|{\bf P}^{\top}{\bf A}{\bf X}{\boldsymbol{\theta}}-{\bf P}^{\top}{\bf Y}\|^{2}$ (16)
@@
Theorem A.1 (woodruff2014sketching, Thm. 23).
Let $\bf P$$^{\top}$ be a sparse subspace embedding with $O(\frac{d^{2}}{\epsilon})$ rows, and let $\boldsymbol{\theta}$$^{*}$ be the solution of Equation (15) and ${\boldsymbol{\theta}}_{C}^{*}$ be the solution of Equation (16). Then, with probability 0.99, it holds that:
$\mathcal{L}_{c}({\boldsymbol{\theta}}_{C}^{*})={\frac{1}{2N}}\|{\bf P}^{\top}{\bf A}{\bf X}{\boldsymbol{\theta}}_{C}^{*}-{\bf P}^{\top}{\bf Y}\|^{2}={\frac{1\pm\epsilon}{2N}}\|{\bf A}{\bf X}{\boldsymbol{\theta}}^{*}-{\bf Y}\|^{2}$ (17)
@@
$\mathcal{L}_{c}({\boldsymbol{\theta}}_{C}^{*})$ $=$ ${\frac{1}{2N}}\|{\bf P}^{\top}{\bf A}{\bf P}{\bf P}^{\top}{\bf X}{\boldsymbol{\theta}}_{C}^{*}-{\bf P}^{\top}{\bf Y}\|^{2}={\frac{1}{2N}}\|{\bf P}^{\top}{\bf A}{\bf X}{\boldsymbol{\theta}}_{C}^{*}-{\bf P}^{\top}{\bf Y}+{\bf R}\|^{2}$
@@
Here we used norm properties and Theorem˜A.1, and $\bf R$} is the residual.
@@
Let $\boldsymbol{\theta}$$^{*}$ be the solution of equation (15) and let ${\boldsymbol{\theta}}_{C}^{*}$ be the solution of equation (16). Denote $\epsilon>0$, and let $\bf P$$^{\top}$ be a sparse subspace embedding with $O(\frac{c^{2}}{\epsilon})$ rows, where $c$ is the size of the feature vectors (number of graph channels), that injects any vector into a subset of indices, which we denote by $C$. According to this subset, denote ${\bf A}=\begin{bmatrix}{\bf A}_{CC}&{\bf A}_{CF}\\ {\bf A}_{FC}&{\bf A}_{FF}\end{bmatrix}$, and ${\bf X}=\begin{bmatrix}{\bf X}_{C}\\ {\bf X}_{F}\end{bmatrix}$. Then if $\sigma=Id$, there exists $\|{\bf R}\|<\|{\bf A}_{CF}{\bf X}_{F}{\boldsymbol{\theta}}_{C}^{*}\|$ such that with probability .99:
$\mathcal{L}_{c}({\boldsymbol{\theta}}_{C}^{*})={\frac{1}{2N}}\|{\bf P}^{\top}{\bf A}{\bf P}{\bf P}^{\top}{\bf X}{\boldsymbol{\theta}}_{C}^{*}-{\bf P}^{\top}{\bf Y}\|^{2}\leq{\frac{1\pm\epsilon}{2N}}\left(\|{\bf A}{\bf X}{\boldsymbol{\theta}}^{*}-{\bf Y}\|^{2}+\|{\bf R}\|^{2}\right)$ (19)
The approximation quality depends on the residual matrix $\bf R$}. For separable graphs, there are no edges that are omitted and therefore ${\bf R}=0$. As the number of edges that are omitted increases, the solution of the coarse-scale problem diverges from the solution of the fine-scale one. While theoretically the residual may become large in certain scenarios, our empirical results show that in practical settings $\bf R$} remains moderate.
@@ Appendix B Qtips and MNIST Datasets
We evaluate our methods on the Qtips dataset by constructing a training set of 100 rods and a test set of 1,000 rods. We employ a 4-layer GCN with 256 hidden channels and compare the performance of our approach across three levels of coarsening. The baseline is trained for 500 epochs, while the multiscale methods are trained for [400, 200, 100] epochs across levels. Results, shown in Table˜6, indicate that the combination of random coarsening and subgraph pooling achieves the best performance, particularly when using $p=3$.
We further test our approach on a graph dataset derived from the MNIST image dataset (lecun1998mnist). The original MNIST images consist of $28\times 28$ grayscale pixels, each depicting a digit from 0 to 9. We convert each image into a graph by treating each pixel as a node and connecting nodes using their spatial positions and a $k$-nearest-neighbor algorithm. The task is to assign each node either to a background class or to the digit class it belongs to. We train a 4-layer GCN with 256 hidden channels on 1,000 graphs and evaluate on 100 test graphs. Results are presented in Table˜7, where we again observe that the combination of random coarsening and subgraph pooling achieves the best performance among the methods evaluated.
@@
Appendix C Coarsening Methods
Additional implementation details of our coarsening procedures are provided. Algorithm˜3 presents the random coarsening algorithm, Algorithm˜4 outlines the Top-$k$ coarsening method, and Algorithm˜5 describes the subgraph generation process.
Algorithm 3 Random Coarsening
# $m$ – coarsining ratio.
Initialize Adj $\bf A$}
for $i=0,\ldots,p-1$ do
  ${\bf A}={\bf A}\times$ Adj
end for
Randomly choose m nodes
Use chosen nodes to calculate the injection matrix $\bf P$}
Use $\bf P$} to calculate ${\bf A}_{r}$ as in Equation (5)
Algorithm 4 Topk Coarsening
# $m$ – coarsining ratio.
Initialize Adj $\bf A$}
for $i=0,\ldots,p-1$ do
  ${\bf A}={\bf A}\times$ Adj
end for
Select $m$ nodes with the largest degrees
Use chosen nodes to calculate the injection matrix $\bf P$}
Use $\bf P$} to calculate ${\bf A}_{r}$ as in Equation (5)
Algorithm 5 Subgraph Coarsening
# $m$ – coarsining ratio.
# $k_{r}$khopk-hop parameter for each level $r$.
Initialize Adj $\bf A$}
for $i=0,\ldots,p-1$ do
  ${\bf A}={\bf A}\times$ Adj
end for
Choose a random node
Generate subgraph using $k_{r}-hops$
Use chosen nodes to calculate the injection matrix $\bf P$}
Use $\bf P$} to calculate ${\bf A}_{r}$ as in Equation (5)
Appendix D Loss Comparison between Multilevel Training and Single-Level Training
To further demonstrate the efficiency of our methods, we show in Figure˜5 a comparison between the loss during training, both using single-level training and multiscale training. We show that both training approaches converge to similar loss after a similar number of training epochs, while as demonstrated in Table˜2, the cost of a single iteration on the coarse graph is lower than on the original fine graph. The plot shows results for the OGBN-Arxiv dataset (ogbn_arxiv), using GCN (gcn) and random pooling.
Table˜8 presents the $\gamma_{r}$ values obtained during multiscale training, computed according to Equation 9. We note that $\gamma_{r}$ increases as the value of $r$ grows.
@@ Appendix D Loss Comparison between Multilevel Training and Single-Level Training
Table 8: Demonstration of $\gamma_{r}$ values, as described in Equation 9.
$\gamma_{r}$ $r=1$ $r=2$ $r=3$
Random 0 0.060 0.076
Topk 0 0.046 0.175
Subgraph 0 0.067 0.170
Appendix E Memory Consumption
We show memory consumptions for OGBN-Arxiv (ogbn_arxiv), as function of coarsening layer and network. The results are summarized in Table˜9, and they show that coarse grids consume less memory than the original fine scale data.
Table 9: Memory consumption across layers and models.
$L=2$ $L=3$ $L=4$
GCN
Fine Grid ($L=1$) 1610.30
  Random 926.62 528.33 332.99
  Topk 994.36 582.51 366.58
  Subgraphs 641.59 491.47 221.85
  Rand. & Sub. 923.39 528.28 333.20
GIN
Fine Grid ($L=1$) 2188.64
  Random 1227.56 719.56 431.26
  Topk 1259.72 739.62 443.28
  Subgraphs 938.48 498.18 224.91
  Rand. & Sub. 1228.66 718.55 431.26
GAT
Fine Grid ($L=1$) 4059.53
  Random 1415.95 588.60 324.92
  Topk 2787.29 1655.82 896.41
  Subgraphs 1108.56 466.31 220.81
  Rand. & Sub. 1423.09 601.40 322.41
Appendix F Cora, CiteSeer and PubMed Datasets
We evaluate our method on the Cora, Citeseer, and PubMed datasets (cora_dataset) using the standard split from cora_split, with 20 nodes per class for training, 500 validation nodes, and 1,000 testing nodes. Cora contains machine learning research papers classified into topics, CiteSeer includes scientific publications from various computer science fields, and PubMed consists of biomedical papers related to diabetes, where nodes represent documents and edges represent citation links. Our experiments cover standard architectures—GCN (gcn), GIN (gin), and GAT (gat)—and are summarized in Table 10, where we achieve performance that is either superior to or on par with the fine-grid baseline, which trains directly on the original graphs.
We explore two pooling strategies—random and Top-k—along with subgraph generation via $k$-hop neighborhoods and a hybrid approach combining both. Each method is tested under two connectivity settings: the original adjacency ($p=1$) and an enhanced version ($p=2$), where we square the adjacency matrix. We report test accuracy across three runs with different random seeds and include standard deviation to assess robustness. Our results indicate that enhanced connectivity improves accuracy in some settings. Table 10 further presents three levels of multiscale training alongside theoretical FLOP analyses, demonstrating the efficiency of our approach. The details of the FLOPs analysis are provided in Section 5, and additional results for two and four levels of multiscale training are provided in Table˜11 and Table˜12. We provide further experimental details in Table˜18, and additional information about the datasets appears in Table˜14.
@@ Appendix F Cora, CiteSeer and PubMed Datasets
  Random(p=2) $2.63\times 10^{2}$ 83.0% (0.5%) $82\times 10^{2} 66.$0% (0.8%) $9.99\times 10^{2}$ 73.8% (0.4%)
@@
  Random(p=1) $2.65\times 10^{2}$ 80.3% (0.2%) $8.11\times 10^{2}$ 72.9% (1.5%) $29\times 10^{2} 77.$5% (1.5%)
@@
Appendix G Additional Transductive Learning Datasets
We further test our methods using additional datasets. We provide results for Flickr flickr and PPI (transductive) facebook_blogcatalog_ppi, showing both timing analysis, in Table˜13, and full multiscale training results in Table˜15. Flickr is a social network where nodes represent images and edges represent mutual user connections or shared tags, focusing on image-based community detection, while PPI (transductive) is a protein-protein interaction network where nodes represent proteins and edges denote biological interactions, with the goal of predicting protein functions within a single connected graph. While pooling does not produce a clearly interpretable or meaningful molecule in the molecular setting, this does not affect the applicability or performance of our methods. Our results show that while achieving comparable accuracy to the original fine grid, our methods show significant timing improvement both in a theoretical manner as well as in FLOPs analysis, which we calculate as explained in Section˜5.
In addition, we test our methods and show results for Facebook, BlogCatalog (facebook_blogcatalog_ppi), DBLP (dblp) and WikiCS (wikics), in Table˜16 and Table˜17. Facebook represents a social network where nodes are users and edges represent friendships, BlogCatalog is a social network of bloggers connected based on shared interests, DBLP is a co-authorship network built from academic publications in computer science, and WikiCS consists of Wikipedia articles about computer science topics connected by hyperlinks. This further illustrates that our methods are not restricted to a specific dataset or network architecture, as our results are consistently comparable to the original training baseline results. Details of our experiments are provided in Table˜18, and we list additional details of the datasets in Table˜14.
We provide additional comparison between our results and sampling-based efficient GNN training methods, which include GraphSAINT (zeng2020graphsaint), FastGCN (chen2018fastgcn), and ClusterGCN (chiang2019cluster). Table˜19 shows a comparison with our methods, for the Flickr (flickr) dataset. We observe that our efficient methods have yielded better results for this dataset in comparison to the above methods.
@@ Appendix G Additional Transductive Learning Datasets
Subgraphs $4.92\times 10^{3}$ 73.16 54.61% (0.8%) $61\times 10^{3} 11$3.90 71.63% (0.3%)
@@
Topk $8.93\times 10^{3}$ 34.47 51.99% (5.1%) $54\times 10^{3} 33.$98 84.92% (0.8%)
@@
Cora cora_dataset 2,708 10,556 1,433 7
Citeseer cora_dataset 3,327 9,104 3,703 6
Pubmed cora_dataset 19,717 88,648 500 3
Flickr flickr 89,250 899,756 500 7
WikiCS wikics 11,701 216,123 300 10
DBLP dblp 17,716 105,734 1,639 4
Facebook facebook_blogcatalog_ppi 4,039 88,234 1,283 193
BlogCatalog facebook_blogcatalog_ppi 5,196 343,486 8,189 6
PPI (transductive) facebook_blogcatalog_ppi 56,944 1,612,348 50 121
@@
Network architecture GCN gcn with 4 layers and 192 hidden channels. GIN gin with 3 layers and 256 hidden channels. GAT gat with 3 layers and 64 hidden channels, using 2 heads
Loss function Negative Log Likelihood Loss except for PPI (transductive) and Facebook facebook_blogcatalog_ppi which use Cross Entropy Loss
Optimizer Adam kingma2014adam
@@
Sub-to-Full strategy Using ego-networks ego_networks. 6 hops on level 2, 4 on level 3, and 2 on level 4
@@
Table 19: Comparisson with results as showes in zeng2020graphsaint, in Micro F1 score
Flickr
GCN 0.492
GraphSAGE 0.501
FastGCN 0.504
S-GCN 0.482
AS-GCN 0.504
ClusterGCN 0.481
GraphSAINT-Node 0.507
GraphSAINT-Edge 0.510
GraphSAINT-RW 0.511
GraphSAINT-MRW 0.510
Ours 0.538
Appendix H Additional Results for Shapenet Dataset
We test out methods on the full Shapenet dataset shapenet using DGCNN dgcnn and demonstrate the results of our multiscale training process in Table 20. We observe that our method is not limited to the ’Airplane’ category of this dataset, and we achieved comparable results across various connectivity values and pooling methods while using a significantly cheaper training process.
To emphasize the efficiency of our methods, we further provide timing analysis of Shapanet ’Airplanes’ data computations. Timing was calculated using GPU computations, and we averaged across 5 epochs after the training loss was stabilized. Results appear in Table˜21 for multiscale training and in Table˜22 for multiscale gradient computations, where we show that our training process was indeed faster than the original process, which only took into consideration the fine grid. We provide additional experimental details in Table˜23.